# A Comprehensive Survey of Deep Transfer Learning for Anomaly Detection in Industrial Time Series: Methods, Applications, and Directions

### v1

@@ Abstract
Automating the monitoring of industrial processes has the potential to enhance efficiency and optimize quality by promptly detecting abnormal events and thus facilitating timely interventions. Deep learning, with its capacity to discern non-trivial patterns within large datasets, plays a pivotal role in this process. Standard deep learning methods are suitable to solve a specific task given a specific type of data. During training, the algorithms demand large volumes of labeled training data. However, due to the dynamic nature of processes and the environment, it is impractical to acquire the needed data for standard deep learning training for every slightly different case anew. Deep transfer learning offers a solution to this problem. By leveraging knowledge from related tasks and accounting for variations in data distributions, this learning framework solves new tasks even with little or no additional labeled data. The approach bypasses the need to retrain a model from scratch for every new setup and dramatically reduces the labeled data requirement. This survey provides an in-depth review of deep transfer learning, examining the problem settings of transfer learning and classifying the prevailing deep transfer learning methods. Moreover, we delve into applying deep transfer learning in the context of a broad spectrum of time series anomaly detection tasks prevalent in primary industrial domains, e.g., manufacturing process monitoring, predictive maintenance, energy management, and infrastructure facility monitoring. We conclude this survey by underlining the challenges and limitations of deep transfer learning in industrial contexts. We also provide practical directions for solution design and implementation for these tasks, leading to specific, actionable suggestions.
Keywords deep transfer learning  $\cdot$ time series analysis  $\cdot$ anomaly detection  $\cdot$ manufacturing process monitoring  $\cdot$ predictive maintenance
@@ 1 Introduction
The fourth industrial revolution — Industry 4.0 (Kagermann et al., 2011), that is characterized by increasing efficiency through the digitization of production, automation, and horizontal integration across companies (Roblek et al., 2016), and the advent of connected cyber-physical systems – referred to as internet of things (Wang et al. 2015; Jeschke et al. 2017; Dalenogare et al. 2018), increases the need for autonomous and intelligent process monitoring. This can be exemplified by the use case of a smart factory in which industrial processes are transformed to be more flexible, intelligent, and dynamic (Kagermann, 2017), or the use case of decentralized energy production with wind and solar (Abir et al., 2021). In these examples, AI-powered anomaly detection integrates the analysis of time series data to detect unusual patterns in the recorded data. By identifying parameters that fall outside a window of normal operation, operators can trigger interventions and adjustments to ensure high product quality and safe operations. To achieve this, physical properties such as pressure or temperature are monitored and analyzed in real-time. Changes in these variables capture drifting and abrupt faults caused by process failures or malfunctions (Park et al., 2020). The production process must adapt quickly to changes in production and the environment to meet the requirements for flexibility and dynamics. Further use cases exist in such diverse areas as manufacturing monitoring including automatic quality control, predictive maintenance of goods and services, infrastructure monitoring of e.g. building energy systems or power plants, digital agriculture, petrochemical process optimization, computer network intrusion detection, or aircraft flight monitoring, to name a few.
@@
Artificial Intelligence, in particular deep learning, provides competent frameworks to automate intelligent monitoring in order to provide valuable assistance to operators and high-level control systems. Leveraging the power of deep learning, formative features of the data – technically referred to as representations (Bengio et al., 2013) – can be captured in a machine-learned model and thereby enable a detailed understanding of variations in standard operations.
@@
However, in non-trivial and non-stationary conditions, the task or the underlying data may change. For example, the monitoring system of a milling machine may in one instance be tasked to detect a blunt tool based on the vibration and in another instance – using the same vibration measurements – to detect insufficient cooling lubricant. Knowledge acquired to solve one task in one setting with a given tool, machined part, and type of machine may be transferred to solve the same or similar task in a setting with a different tool, machined part, or type of machine. Slowly changing conditions (drifts), abrupt mode changes (for instance due to tool change), and new tasks (such as the detection of another failure mode) may require adjustments to the machine-learned model. In these cases, it is desirable to adjust the analysis model without retraining from scratch, as it is costly or impractical to acquire sufficient training data to learn the full manifold (Maschler et al., 2022).
@@
Transfer learning is a machine learning framework to achieve this (Maschler and Weyrich, 2021; Zhuang et al., 2021; Pan and Yang, 2010; Tan et al., 2018; Yosinski et al., 2014). As depicted in Figure 1, data and algorithms from a related task may be leveraged in a new one. By accounting for changes in data distributions and tasks, or leveraging existing models, knowledge learned from related tasks can be used to improve performance on new tasks instead of retraining a model for each individual application from scratch. This transfer-learning-boosted modeling forms the basis for identifying anomalies that deviate from established patterns in a non-trivial manner without full re-training.
@@
Figure 1: Transfer learning is useful when changes in production take place and sufficient data for full retraining is not available as shown here for a hypothetical production of two types of gears. In the production of gear A, a lot of data is available to train a deep learning model that helps improve production. In the production of gear B, data availability is more limited, and the traditionally trained deep learning model fails to improve production. With suitable methods, however, data and algorithms acquired during production of gear A can be leveraged to support improving the production of gear B, because the data and tasks in the production of both gears are related.
@@
Deep transfer learning (Tan et al., 2018; Yu et al., 2022) extends the transfer learning paradigm by leveraging deep learning models. In industrial contexts, it ensures optimal production even as production conditions shift. This dynamic adaptability is key in maintaining the effectiveness of anomaly detection systems in the dynamic environment that characterizes industrial applications including the broad categories of manufacturing process monitoring, predictive maintenance, energy management, and infrastructure facility monitoring as detailed in Section 4.
@@
In this survey, we review the foundations of deep transfer learning to equip the reader with working knowledge of the main principles and intuition for ideas. Further, we provide a comprehensive overview of the current state of the art of deep transfer learning approaches for time series anomaly detection for industrial applications. Our main contribution is a systematic review of research work on real-world industrial applications. For these, we discuss potential, challenges, and limitations and give directions for future work and potential. The paper is organized as follows: First, we introduce a taxonomy of transfer learning problem settings and further categorization of deep transfer learning approaches (Section 2). Then, we describe the task of anomaly detection in time series (Section 3) in selected industrial applications (Section 4). To conclude, we discuss current challenges, limitations, and future research directions (Sections 56) in the field.
@@ 2 Deep transfer learning
Transfer learning in a deep learning setting aims to increase the efficiency, performance, and generalization of deep learning models by transferring knowledge from one data set and task to a new one. This eliminates the need to train a deep learning model from scratch, which in turn reduces the amount of necessary data and compute required to solve a new task or new data domain. In either case, knowledge is transferred from a source to a target domain, as defined below. The transfer learning problem settings can be categorized as inductive or transductive transfer depending on the data and task conditions, while we categorize deep learning-based transfer learning approaches into instance transfer, parameter transfer, mapping transfer and domain-adversarial transfer. We illustrate them by using two intuitive examples in Figure 2, with more details being elaborated in the following sections.
@@
Figure 2: Taxonomy of transfer learning problem settings (left; see Section 2.1 for the definition of terms) and corresponding examples using deep transfer learning approaches (right). On the left, we classify transfer learning problems as inductive or transductive transfer settings, mainly depending on the label availability and task similarities. Correspondingly, we provide two examples using deep transfer learning methods: In the inductive transfer setting, we collect time series data from screw production and wrench production. Labeled screw data (A1) is used to detect collective anomalies (a set of data points behaving differently compared to the entire time series (Choi et al., 2021; Chalapathy and Chawla, 2019), further explained in Section 3). Then, parameter transfer (Section 2.2.2) is applied to transfer knowledge by fine-tuning the pre-trained model from labeled screw data to detect point anomalies (further explained in Section 3) on labeled wrench data (A2). For the transductive transfer setting in the lower panel, we present a different situation for contextual anomaly detection (further explained in Section 3). In this case, we have two datasets, B1 and B2, analyzed using the same model. However, the data in B2 significantly differs in appearance from the data in B1. To address this issue, instance transfer (further explained in Section 2.2.1) is used. Through this learning process, the data in B2 is transformed in a way that makes it compatible with the model which has been trained exclusively on data from B1. Transfer learning, in this case, is thus achieved by adapting the data to fit the model through domain adaptation rather than adjusting the model to fit new data.
2.1 Transfer learning problem definition
@@
Domain $\mathcal{D}$ includes the domain feature space $\mathcal{X}$ and marginal data distribution $P(X)$ as $\mathcal{D}=\{\mathcal{X},P(X)\}$, where $X$ is the domain data, $X=\left\{x_{1},\ldots,x_{n}\right\}\in\mathcal{X}$. Similarly, a learning task is defined as $\mathcal{T}=\{\mathcal{Y},f_{\mathcal{T}}(\cdot)\}$, where $\mathcal{Y}$ denotes the task space and usually represents class label. For anomaly detection tasks, $\mathcal{Y}$ is the set of the two classes “normal” and “abnormal”. The function $f_{\mathcal{T}}(\cdot)$ can be used to predict the corresponding label of a new instance $x_{i}$. The objective predictive function $f_{\mathcal{T}}(\cdot)$ can be learned from domain data and can be interpreted as a form of conditional probability. Thus, the learning task can be rewritten as $\mathcal{T}=\{\mathcal{Y},P(Y|X)\}$, where $P(Y|X)$ is used as a likelihood measure to determine how well a given data set $X$ fits with a corresponding class label set $Y$.
@@
We largely follow the definition of transfer learning by Pan and Yang (2010) and Zhuang et al. (2021). Given a source domain $\mathcal{D}_{S}$ and learning task $\mathcal{T}_{S}$, as well as a target domain $\mathcal{D}_{T}$ and learning task $\mathcal{T}_{T}$, transfer learning aims to improve the performance of the predictive function $f_{\mathcal{T}}(\cdot)$ in $\mathcal{D}_{T}$ by transferring knowledge from $\mathcal{D}_{S}$ and $\mathcal{T}_{S}$, where $\mathcal{D}_{S}\neq\mathcal{D}_{T}$ and/or $\mathcal{T}_{S}\neq\mathcal{T}_{T}$. Usually, the size of source dataset is much smaller than target dataset.
@@
This definition of transfer learning can be broadened, i.e., the target task can profit from multiple source domains. Transfer learning is thus the idea of making the best use of related source domains to solve new tasks. In contrast, traditional machine learning (ML) methods learn each task separately from scratch, and each respective model can only be applied to the corresponding task.
@@
We define a taxonomy of transfer learning problems settings as shown in Figure 2 mainly depending on the label availability in the two domains to be easily applicable to the requirements of a case at hand (compare different definitions for other purposes in the literature Pan and Yang (2010); Tan et al. (2018); Zhuang et al. (2021)).
@@
We differentiate it into inductive and transductive transfer learning (Pan and Yang, 2010) [^1] In this survey, we do not consider unsupervised learning scenarios since either source labels or target labels are provided for most industrial applications.. Inductive transfer learning is applied when the target task is different from the source task, i.e., $\mathcal{T}_{S}\neq\mathcal{T}_{T}$ (meaning that $\{\mathcal{Y_{S}}\neq\mathcal{Y_{T}}\}$ or $\{P(Y_{S}|X_{S})\neq P(Y_{T}|X_{T})\}$). The conditional probability distribution is induced with labeled training data in the target domain (Torrey and Shavlik, 2009). A corresponding example is illustrated as Scenario A in Figure 2, where the learning tasks are different and the goal of transfer learning is to recognize point anomaly from the collective anomaly task. Related areas of inductive transfer learning are multi-task learning (Caruana, 1997; Ruder, 2017) and sequential learning, depending on whether tasks are learned simultaneously or sequentially.
@@
Transductive transfer learning is applied when the source and target tasks are the same, while the source and target domain are different, i.e., $\mathcal{T}_{S}=\mathcal{T}_{T}$ and $\mathcal{D}_{S}\neq\mathcal{D}_{T}$ (meaning that $\{\mathcal{X_{S}}\neq\mathcal{X_{T}}\}$ or $\{P(X_{S})\neq P(X_{T})\}$. A subcategory is domain adaptation (Kouw and Loog, 2019) when the feature space of source and target data are the same but the corresponding marginal distributions are different (i.e., $\{\mathcal{X_{S}=X_{T}}\}$ and $\{P(X_{S})\neq P(X_{T})\}$). Scenario B in Figure 2 is an example of transductive transfer learning where the learning tasks are identical, and the goal of transfer learning is to recognize contextual anomalies in an unlabelled data set.
Other learning paradigms closely related to transfer learning are listed below:
Multi-task learning is a machine learning technique where a single model is trained on multiple tasks simultaneously. The idea is to improve the performance of the model by learning a shared representation that captures the features between all tasks.
Continuous learning (Parisi et al., 2019) is a learning process where the model continuously learns new tasks from previous tasks over time without forgetting how to solve previous tasks. To some extent, continuous learning can be seen as a sequential transfer learning process, with the constraint to preserve the performance on the previous tasks which leads to an accumulation of knowledge over time.
Few-shot learning (Wang et al., 2020) is a type of machine learning where a model can learn and perform well on a new task with only a limited number of labeled samples. In extreme cases, the model can learn with one label (Fei-Fei et al., 2006) and without any label (Lampert et al., 2009). While transfer learning usually involves reusing the model from relevant tasks and continuing training on the target dataset.
Meta-learning (Finn et al., 2017; Hospedales et al., 2022) is a machine learning technique that focuses on the learning process. It is known as “learning to learn”. For meta-learning, models are trained on a different set of tasks instead of a set of data in the traditional machine learning setting. In this sense, meta-learning can be seen as a form of transfer learning because it involves transferring knowledge from task to task.
Knowledge distillation (Gou et al., 2021) effectively learns a small model trained to mimic the behavior of a larger, more complex model. The knowledge learned by the larger model can be transferred to the smaller model, which can then be used for the target task, e.g., on a less powerful edge device.
Self-supervised learning (Liu et al., 2023; Bai et al., 2021) involves training a model to predict some aspect of the input data without any external supervision. The learned representations can be used for various downstream tasks, including those that involve transferring knowledge from one domain to another.

### v2

@@ Abstract
Automating the monitoring of industrial processes has the potential to enhance efficiency and optimize quality by promptly detecting abnormal events and thus facilitating timely interventions. Deep learning, with its capacity to discern non-trivial patterns within large datasets, plays a pivotal role in this process. Standard deep learning methods are suitable to solve a specific task given a specific type of data. During training, deep learning demands large volumes of labeled data. However, due to the dynamic nature of the industrial processes and environment, it is impractical to acquire large-scale labeled data for standard deep learning training for every slightly different case anew. Deep transfer learning offers a solution to this problem. By leveraging knowledge from related tasks and accounting for variations in data distributions, the transfer learning framework solves new tasks with little or even no additional labeled data. The approach bypasses the need to retrain a model from scratch for every new setup and dramatically reduces the labeled data requirement. This survey first provides an in-depth review of deep transfer learning, examining the problem settings of transfer learning and classifying the prevailing deep transfer learning methods. Moreover, we delve into applications of deep transfer learning in the context of a broad spectrum of time series anomaly detection tasks prevalent in primary industrial domains, e.g., manufacturing process monitoring, predictive maintenance, energy management, and infrastructure facility monitoring. We discuss the challenges and limitations of deep transfer learning in industrial contexts and conclude the survey with practical directions and actionable suggestions to address the need to leverage diverse time series data for anomaly detection in an increasingly dynamic production environment.
@@ 1 Introduction
The fourth industrial revolution – Industry 4.0 [1], that is characterized by increasing efficiency through the digitization of production, automation, and horizontal integration across companies [2], and the advent of connected cyber-physical systems – referred to as internet of things [3, 4, 5], increases the need for autonomous and intelligent process monitoring. This can be exemplified by the use case of a smart factory in which industrial processes are transformed to be more flexible, intelligent, and dynamic [6], or the use case of decentralized energy production with wind and solar [7]. In these examples, AI-powered anomaly detection integrates the analysis of time series data to detect unusual patterns in the recorded data. To achieve this, a deep learning architecture is modeled to capture indicators of normal and abnormal operation. The learning process involves the analysis of historic time series sensor data of normal and possibly abnormal operations. This data is for example used for representation- or reconstruction-based learning. After training, the deep learning model represents or reconstructs normal data in a certain way. The model is designed in a way that abnormal data–because it is different–is either represented differently from the normal data or reconstructed poorly and thus recognized as an anomaly. By identifying operational parameters that fall outside a window of normal interval, operators can trigger interventions and adjustments to ensure high product quality and safe operations. To achieve this, physical properties such as pressure or temperature are monitored and analyzed in real-time applications. Changes in these variables capture drifting and abrupt faults caused by process failures or malfunctions [8]. The production process must adapt quickly to changes in production and the environment to meet the requirements for flexibility and dynamics. Further use cases exist in a wide range of diverse fields, such as manufacturing monitoring including automatic quality control [9, 10], predictive maintenance of goods and services [11, 12, 13, 14, 15, 16], infrastructure monitoring of building systems [17, 18] and power plant [19], digital agriculture [20], petrochemical process optimization [21], computer network intrusion detection [22], or aircraft flight monitoring [23], to name a few.
@@
Artificial Intelligence, particularly deep learning, provides competent frameworks with underlying deep neural networks to automate intelligent monitoring and provide valuable assistance to operators and high-level control systems. Leveraging the power of deep learning, informative features of the data – technically referred to as representations [24] – can be captured in a machine-learned model and thereby enable a detailed understanding of variations in standard operations.
@@
However, the task or underlying data may change under non-trivial and non-stationary conditions. For instance, the monitoring system of a milling machine may be assigned the task of identifying a blunt tool based on vibration in one scenario, and in a different scenario, it may utilize the same vibration measurements to detect insufficient cooling lubricant. Knowledge acquired to solve one task in one setting with a given tool, machined part, and type of machine may be transferred to solve the same or similar task in another setting with a different tool, machined part, or type of machine. Slowly changing conditions (drifts), abrupt mode changes (for instance, due to tool change), and new tasks (such as the detection of another failure mode) may require adjustments to the deep learning model. In these cases, it is desirable to adjust the analysis model without retraining from scratch, as it is costly or impractical to acquire sufficient training data to learn the full manifold [25].
@@
Transfer learning is a machine learning framework to achieve this [26, 27, 28, 29, 30]. As depicted in Fig. 1, data and algorithms from one task may be leveraged in a new related one. By accounting for changes in data distributions and tasks or leveraging existing models, knowledge learned from related tasks can be used to improve performance on new tasks instead of retraining a model for each individual application from scratch. This transfer-learning-boosted modeling forms the basis for identifying anomalies that deviate from established patterns in a non-trivial manner without full re-training.
@@
Figure 1: Transfer learning is useful when changes in production take place and sufficient data for full retraining is not available as shown here for a hypothetical production of two types of gears. In the production of gear A, a lot of data is available to train a deep learning model that helps improve production. In the production of gear B, data is more limited, and the traditionally trained deep learning model fails to improve production. With suitable transfer learning methods, however, data and algorithms acquired during the production of gear A can be leveraged to support improving the production of gear B because the data and tasks in the production of both gears are related.
@@
Deep transfer learning [29, 31] extends the transfer learning paradigm by leveraging deep learning. In industrial contexts, it ensures optimal production even as production conditions shift. This dynamic adaptability is key in maintaining the effectiveness of anomaly detection systems in the dynamic environment that characterizes industrial applications including the broad categories of manufacturing process monitoring, predictive maintenance, energy management, and infrastructure facility monitoring as detailed in Section 4.
This survey is a non-systematic yet application-oriented review with a narrow focus on deep transfer learning for anomaly detection in time series in the industry. Our main contributions are as follows:
  • We categorize transfer learning problem settings and then systematically summarize deep transfer learning approaches into four categories. With the foundations of deep transfer learning, we equip the reader with a working knowledge of the main principles and intuitions.
  • We analyze the recent literature and provide a comprehensive overview of the current state of the art of deep transfer learning approaches for time series anomaly detection for main industrial applications.
  • We discuss potential challenges and limitations and then give directions for future work with actionable recommendations for AI practitioners and decision-makers.
To our knowledge, this is the first survey of deep transfer learning in the narrow context of industrial time series anomaly detection. The review describes the underlying methodological principles and methods within a generic taxonomy and discusses practical implications for AI practitioners to make informed decisions. We cover multiple areas of application, including manufacturing monitoring, maintenance prediction, and infrastructure monitoring.
@@
The rest of the paper is organized as follows. First, we provide an overview of transfer learning by introducing a taxonomy of transfer learning problem settings and further categorizing deep transfer learning approaches (Section 2). Then, we describe the task of anomaly detection in time series (Section 3) in selected industrial applications (Section 4). To conclude, we discuss current challenges, limitations, and future research directions (Sections 56) in the field.
1.2 Survey methodology
We seek to identify application-oriented peer-reviewed literature in the intersection of transfer learning as the learning framework, time series as the data domain, and anomaly detection as the task (Fig. 2). To execute the selection process of literature, we search related terms on Google Scholar, Scopus, Elsevier, and IEEE databases. Based on the title, we pick those papers that may fit the narrow topic into a pre-selection list. Eventually, we included publications matching the topic according to the abstract and screening of the content.
Along the reviewed topical papers, we include contextually relevant papers such as deep learning approaches that are agnostic to data types and tasks. For deep transfer learning in general, we searched the keywords “transfer learning” and “deep transfer learning”. Specifically, we focus more on deep transfer learning approaches. Then, we switch to the application-oriented cases where deep transfer learning is applied to tackle time series anomaly detection in the main industrial applications. To achieve this, we search queries like “deep transfer learning for time series anomaly detection” and “deep transfer learning for predictive maintenance”. After searching in the database, we carefully check and screen out the most relevant literature based on the following inclusion/exclusion criteria: (1) We only include the applications that utilize deep transfer learning approaches, instead of traditional transfer learning; (2) We only include publications after 2013; (3) We cover all three main topics in Fig. 2 (highlighted with cycles), but we specifically focus on the intersection of the three aforementioned domains. After carefully screening out, we select 45 papers for deep transfer learning in general and 37 papers for deep transfer learning for anomaly detection in industrial time series.
Fig. 3 illustrates the taxonomy in this survey to categorize reviewed studies based on different aspects, including deep transfer learning, time series anomaly detection, industrial applications, current challenges, and future directions.
Figure 2: Venn diagram of this survey’s focus on the intersection of transfer learning, anomaly detection, and time series analysis.
Figure 3: A generic taxonomy in this paper to analyze deep transfer learning for industrial time series anomaly detection.
@@ 2 Deep transfer learning
2.1 Overview of the field
[…]
@@ 2.1 Overview of the field
Figure 4: Taxonomy of transfer learning problem settings (left; see Section 2.2 for the definition of terms) and corresponding examples using deep transfer learning approaches (right). On the left, we classify transfer learning problems as inductive or transductive transfer settings Correspondingly, we provide two examples using deep transfer learning methods: In the inductive transfer setting, we collect time series data from screw production and wrench production. Labeled screw data (A1) is used to detect collective anomalies (a set of data points behaving differently compared to the entire time series [32, 33], further explained in Section 3). Then, parameter transfer (Section 2.3.2) is applied to transfer knowledge by fine-tuning the pre-trained model from labeled screw data to detect point anomalies (further explained in Section 3) on labeled wrench data (A2). For the transductive transfer setting in the lower panel, we present a different situation for contextual anomaly detection (further explained in Section 3). In this case, we have two datasets, B1 and B2, analyzed using the same model. However, the data in B2 significantly differs in appearance from the data in B1. To address this problem, instance transfer (further explained in Section 2.3.1) is used. Through this learning process, the data in B2 is transformed in a way that makes it compatible with the model that has been trained exclusively on data from B1. Transfer learning, in this case, is thus achieved by adapting the data to fit the model through domain adaptation rather than adjusting the model to fit new data.
2.2 Formal description of deep transfer learning
@@
Domain $\mathcal{D}$ includes the domain feature space $\mathcal{X}$ and marginal data distribution $P(X)$ as $\mathcal{D}=\{\mathcal{X},P(X)\}$, where $X$ is the domain data, $X=\left\{x_{1},\ldots,x_{n}\right\}\in\mathcal{X}$. Similarly, a learning task is defined as $\mathcal{T}=\{\mathcal{Y},f_{\mathcal{T}}(\cdot)\}$, where $\mathcal{Y}$ denotes the task space and usually represents class label. For anomaly detection tasks, $\mathcal{Y}$ is the set of the two classes “normal” and “abnormal”. The function $f_{\mathcal{T}}(\cdot)$ can be used to predict the corresponding label of a new instance $x_{i}$. The objective predictive function $f_{\mathcal{T}}(\cdot)$ learned from domain data can be interpreted as a form of conditional probability. Thus, the learning task can be rewritten as $\mathcal{T}=\{\mathcal{Y},P(Y|X)\}$, where $P(Y|X)$ is used as a likelihood measure to determine how well a given data set $X$ fits with a corresponding class label set $Y$.
@@
In the surveyed literature on transfer learning for anomaly detection in industrial applications, transfer learning methods from other fields, such as computer vision and natural language processing, were adopted. We therefore use a generic classification scheme for transfer learning methods. We largely follow the definition of transfer learning in literature [27, 28]. Given a source domain $\mathcal{D}_{S}$ and learning task $\mathcal{T}_{S}$, as well as a target domain $\mathcal{D}_{T}$ and learning task $\mathcal{T}_{T}$, transfer learning aims to improve the performance of the predictive function $f_{\mathcal{T}}(\cdot)$ in $\mathcal{D}_{T}$ by transferring knowledge from $\mathcal{D}_{S}$ and $\mathcal{T}_{S}$, where $\mathcal{D}_{S}\neq\mathcal{D}_{T}$ and/or $\mathcal{T}_{S}\neq\mathcal{T}_{T}$. Usually, the size of source dataset is much smaller than target dataset.
@@
This definition of transfer learning can be broadened, i.e., the target task can benefit from multiple source domains. Transfer learning is thus the idea of making the best use of related source domains to solve new tasks. In contrast, traditional machine learning (ML) methods learn each task separately from scratch, and each respective model can only be applied to the corresponding task.
@@
We define a taxonomy of transfer learning problem settings as shown in Fig. 4 mainly depending on the label availability in the two domains to be easily applicable to the requirements of a case at hand (compare different definitions for other purposes in the literature [27][29]).
@@
We differentiate it into inductive and transductive transfer learning [28]. Inductive transfer learning is applied when the target task is different from the source task, i.e., $\mathcal{T}_{S}\neq\mathcal{T}_{T}$ (meaning that $\{\mathcal{Y_{S}}\neq\mathcal{Y_{T}}\}$ or $\{P(Y_{S}|X_{S})\neq P(Y_{T}|X_{T})\}$). The conditional probability distribution is induced with labeled training data in the target domain [34]. A corresponding example is illustrated as Scenario A in Fig. 4, where the learning tasks are different and the goal of transfer learning is to recognize point anomaly from the collective anomaly task. Related areas of inductive transfer learning are multi-task learning [35, 36] and sequential learning, depending on whether tasks are learned simultaneously or sequentially.
@@
Transductive transfer learning is applied when the source and target tasks are the same, while the source and target domain are different, i.e., $\mathcal{T}_{S}=\mathcal{T}_{T}$ and $\mathcal{D}_{S}\neq\mathcal{D}_{T}$ (meaning that $\{\mathcal{X_{S}}\neq\mathcal{X_{T}}\}$ or $\{P(X_{S})\neq P(X_{T})\}$. A subcategory is domain adaptation [37] when the feature space of source and target data are the same but the corresponding marginal distributions are different (i.e., $\{\mathcal{X_{S}=X_{T}}\}$ and $\{P(X_{S})\neq P(X_{T})\}$). Scenario B in Fig. 4 is an example of transductive transfer learning where the learning tasks are identical, and the goal of transfer learning is to recognize contextual anomalies in an unlabelled data set.